OpenStudy (anonymous):

Which of the following equations describe the line shown below? CORDS (-1,6) and (2,-6) All that apply. A. y - 6 = -4(x + 1) B. y - 1 = -4(x + 6) C. y - 2 = -4(x + 6) D. y = -4x + 2 E. y - 6 = -4(x + 2) F. y = -4x + 3

OpenStudy (anonymous):

C. y - 2 = -4(x + 6)

OpenStudy (anonymous):

is that the only one that applies ?

OpenStudy (anonymous):

ya!

OpenStudy (anonymous):

there is one more atleast.

OpenStudy (anonymous):

D. y = -4x + 2

OpenStudy (anonymous):

i guess to use this equation: (y-y1)/(y2-y1)=(x-x1)/x2-x1) where (x1,y1)=(-1,6) and (x2,y2)=(2,-6)

OpenStudy (anonymous):

Answers were A , D.

OpenStudy (paxpolaris):

yes